# Effects of rotation and anisotropy on the properties of type-II holographic superconductors

### v1

@@ 4 The AC conductivity
The AC conductivity is an important quantity characterizing a field theory superconductor as the response of the system to electric perturbations. In our holographic setup, this can be studied by introducing a gauge field perturbation within the bulk dynamics and working the linearized Maxwell equations. Hence, considering the bulk perturbation along the $x$ component of the gauge field and assuming a time dependence of the form $\delta A_{x}=A_{x}(u)e^{-i\omega t}$, from (6) we arrive at the following equation
@@
We solve this equation by imposing an ingoing wave condition at the horizon
@@
while the asymptotic behavior at the boundary has the form
@@
where the standard holographic dictionary allows $A_{x}^{(0)}$ to be interpreted as the source of a conjugated current $J_{x}$, while $A_{x}^{(z+1)}$ corresponds to its expectation value, $\expectationvalue{J_{x}}\propto A_{x}^{(z+1)}$. In this sense, from the Ohm law, we obtain an expression for the AC conductivity
@@
In the remainder of this section, we are interested in investigating the frequency and angular momentum dependence of the conductivity. For this purpose, we conduct an analytical study of the Maxwell perturbation (24) by focussing on the near-boundary limit of the anisotropic background. This is achieved by considering $f\approx 1$ and $\psi\approx\expectationvalue{\mathcal{O}}u^{\Delta}/\sqrt{2}$, yielding
@@
The analytical solutions of this equation are only possible when considering particular values of the parameters $\Delta$ and $z$. In the subsequent pages, we illustrate some examples in which this can be done. In addition, we supplement our analytical computations with numerical ones by solving the full equation (24) subject to the boundary conditions (25) and (26).
4.1 Case $\Delta=1$ and $z=1$
In this case, the equation (28) has two independent solutions that can be expressed in terms of the Bessel functions of the first $J_{1}$ and second kind $Y_{1}$
@@ 4.1 Case $\Delta=1$ and $z=1$
Before proceeding, we find convenient the identification
$\dfrac{c_{0}}{c_{1}}=\dfrac{i}{\pi}\left[\pi H(\omega-\omega_{g})+(1-2\gamma_{E})i+\dfrac{2\omega}{h^{2}(\omega)}\left(\dfrac{950}{\omega}-5\right)+\frac{1}{100}\,i\ln\left(\dfrac{h^{2}(\omega)}{4}\right)-\dfrac{\pi}{2}\right],$ (31)
@@
wherein $H(\omega-\omega_{g})$ is the Heaviside step function and $\gamma_{E}$ is the Euler constant. Thence, the asymptotic analysis of the solution (29), in conjunction with the relation (27), provides the following result for the conductivity
$\sigma(\omega)=-\frac{{h^{2}(\omega)}}{200\omega}\left(201\pi H(\omega-\omega_{g})+101\,i\ln\left(\dfrac{h^{2}(\omega)}{4}\right)-500\,i\,\dfrac{\omega}{h^{2}(\omega)}\left(\frac{190}{\omega}-1\right)\right).$ (32)
@@
A closer inspection of this expression reveals the structure of real and imaginary parts of the conductivity, giving rise to
$\text{Re}[\sigma]=\begin{cases}\quad\infty\qquad&\quad\omega=0,\\ \quad 0\qquad&0<\omega\leq\omega_{g},\\ -\dfrac{\pi}{2}\dfrac{h^{2}(\omega)}{\omega}&\quad\omega>\omega_{g},\\ \end{cases}$ (33)
$\text{Im}[\sigma]=\begin{cases}-\dfrac{101}{200}\dfrac{{h^{2}(\omega)}}{\omega}\ln\left(\dfrac{h^{2}(\omega)}{4}\right)+\dfrac{1}{2}\left(\dfrac{950}{\omega}-5\right)\qquad&0<\omega<\omega_{g},\\ -\dfrac{101}{200}\dfrac{{h^{2}(\omega)}}{\,\omega}\ln\left(\dfrac{\absolutevalue{h^{2}(\omega)}}{4}\right)+\dfrac{1}{2}\left(\dfrac{950}{\omega}-5\right)\qquad&\quad\omega>\omega_{g}.\\ \end{cases}$ (34)
@@
Here, $\omega_{g}=\expectationvalue{\mathcal{O}}/\ell^{2}$ denotes a gap frequency determined by the function $h(\omega)$ below which the real part of the conductivity drops to zero, except at $\omega=0$, where a delta function appears.
@@
In Figure 2, we show the behavior of the real and imaginary parts of the conductivity defined by these expressions. The observed behavior exhibits recurring characteristics found in previous holographic models for superconductors. Concerning the real part of the conductivity, the delta function at $\omega=0$ suggests the presence of a DC current and the onset of the superconducting state. Moreover, the existence of a gap frequency $\omega_{g}$, which is proportional to the order parameter $\expectationvalue{\mathcal{O}}$, is also a common property in holographic superconductors. However, it has been argued that holographic superconductors are not gapped. Instead, the apparent gap is the result of decoupling the matter sector from the gravity, a consequence of employing the probe limit method. For an in-depth exploration of this matter, we refer the reader to Hartnoll:2009sz ; Horowitz:2009ij ; Horowitz:2010gk . For frequencies above the gap frequency $\omega>\omega_{g}$, the behavior is governed by the monotonic increasing function $\text{Re}[\sigma]\sim-h^{2}(\omega)/\omega$, following a linear tendency $\text{Re}[\sigma]\sim\omega$ at sufficiently higher frequencies.
@@
Turning our attention to the imaginary component, we observe the presence of a minimum occurring near $\omega_{g}$. Below this minimum, at lower frequencies, the abrupt increasing behavior is defined by $\text{Im}[\sigma]\sim 1/\omega$. On the other hand, at sufficiently higher frequencies $\omega\gg\omega_{g}$, the behavior is dictated by $\text{Im}[\sigma]\sim\omega\ln\left(\ell^{2}\omega/2r_{h}\right)$. All these behaviors are contained in Figure 2, where we highlight the remarkable consistency between our analytical predictions (solid lines) and the numerical computations (dashed lines).
@@
Lastly, it is important to note that we have not addressed the rotation of the background in the previous scenario. The reason for this is that neither of the expressions, (33) or (34), possess information on the rotation parameter $a$. This observation results evident from the effective equation (28), where the absence of anisotropy ($z=1$) leads to the cancellation of terms associated with rotation, rendering no residual information on this parameter. Consequently, we infer that anisotropy ($z\neq 1$) is an essential ingredient for discerning the impact of rotation on the AC conductivity, at least in the case that concerns our holographic model.
@@
Figure 2: Sketches for the real (a) and imaginary (b) components of the conductivity as a function of the frequency. Note the excellent agreement between the analytical (solid lines) and the numerical results (dashed lines). These plots were generated at $T\approx 0.93\,T_{c}$ with $\expectationvalue{\mathcal{O}}\approx 8.344$, considering parameters $\ell=1$ and $r_{h}=1$.
4.2 Case $\Delta=1$ and $z=2$
@@
In this scenario, the solution can be written in terms of the confluent hypergeometric functions of the first kind, $M$, and the second kind, $U$,
@@
As in the preceding case, we can obtain an analytic expression for the conductivity using the relation (27). To accomplish this, we find it convenient to make the identification
@@
$\mathcal{F}(a)$ $=\dfrac{448}{5}\left(\dfrac{a}{\ell}\right)^{6}-\dfrac{15147}{100}\left(\dfrac{a}{\ell}\right)^{5}+\dfrac{11171}{100}\left(\dfrac{a}{\ell}\right)^{4}-\dfrac{125}{4}\left(\dfrac{a}{\ell}\right)^{3}+\dfrac{337}{100}\left(\dfrac{a}{\ell}\right)^{2}$ (38a)
$\quad-\dfrac{11}{250}\left(\dfrac{a}{\ell}\right)+\dfrac{1601}{600},$
$\mathcal{G}(a,\omega)$ $=\ell\left[9543\,\left(\dfrac{a}{\ell}\right)^{2}-3493\,\left(\dfrac{a}{\ell}\right)+379+\left(10^{6}\left(\dfrac{a}{\ell}\right)^{2}+5\right)\ell\omega+\frac{6\times 10^{7}}{\ell\omega}\right].$ (38b)
Altogether, with the aid of the above definitions, we arrive at the following expression for the conductivity
@@
In contrast with the previous case, the conductivity in this scenario contains an explicit dependency on the rotation parameter $a$. This makes reasonable the questioning of how the rotation influences the AC conductivity, as shown in Figure 3. To address this question, let us first examine the case devoid of rotation, where $a=0$. This scenario is represented by the black curves in Figure 3. In this case, in conjunction with (39), the real part of the conductivity also has a delta function at $\omega=0$. This arises from the behavior of the denominator |$\absolutevalue{\Gamma(-\frac{1}{4}+i\,g(\omega,a))}^{2}$, which goes to zero more rapidly than it does the function in the numerator when $\omega\to 0$. In addition, Figure 3(a) shows an apparent frequency gap in the low-frequency region. We will come back to this matter further down; for now, let us just denote the width of this gap as $\omega_{g}$. For frequencies beyond this gap, $\omega>\omega_{g}$, the function $\omega^{1/2}$ dominates the behavior of $\text{Re}[\sigma]$, establishing a monotonically increasing asymptotic trend growing slower than the isotropic case sketched in Figure 2(a).
@@
When turning on the rotation ($a\neq 0$), the asymptotic behavior of $\text{Re}[\sigma]$ undergoes a significant transformation. In this scenario, the Dirac delta function at $\omega=0$ persists, accompanied by the emergence of a narrow gap at lower frequencies. This discernible gap can be defined by noting that the values assumed by the function $\text{Re}[\sigma]$ within the interval $0<\omega\leq\omega_{g}$ are considerably smaller compared to those beyond this range, $\omega>\omega_{g}$. Consequently, within $0<\omega\leq\omega_{g}$, we can effectively treat $\text{Re}[\sigma]$ as zero. With this assumption, we can estimate the extent of the gap as a function of the rotation parameter, yielding
$\omega_{g}=e^{-(\pi a/2\ell)}\left[\dfrac{18}{5}\quantity(\dfrac{a}{\ell})^{4}-\frac{1167}{100}\quantity(\dfrac{a}{\ell})^{3}+\dfrac{67}{4}\quantity(\dfrac{a}{\ell})^{2}+8\quantity(\dfrac{a}{\ell})+6\right],$ (40)
wherein we observe a reduction of the gap driven by increasing the rotation. Also, this equation shows that the gap persists even in the absence of rotation, as expected from our earlier discussion.
@@
Figure 3: Portraits of the real (a) and imaginary (b) components of the conductivity as a function of the frequency. Again, our analytical computations (solid lines) show an excellent agreement with numerical ones (dashed lines). These plots were generated at $T\approx 0.99994\,T_{c}$ with $\expectationvalue{\mathcal{O}}\approx 7.780$, considering $\ell=1$ and $r_{h}=1$.
It is interesting to notice that, when $\omega>\omega_{g}$, and for a specific nontrivial value of $a$, the curve exhibits a peak, beyond which it undergoes an exponential fall. In this domain, at sufficiently high frequencies, the term $a^{2}\omega/4\ell\Xi$ dominates with respect to $\expectationvalue{\mathcal{O}}/\ell\omega$ contained in $g(\omega,a)$, giving rise to an asymptotic behavior of the form $\text{Re}[\sigma]\sim\omega^{2}\exp\left(-2\pi a^{2}\omega/4\ell\Xi\right)$, which is now characterized by containing a proper dependence on the rotation parameter. Therefore, the combined effect of rotation and anisotropy in the background is accountable for inducing the peak and vanishing exponential behavior. This is a noteworthy characteristic in our model, as it resembles the behavior observed in field theory superconductor models as well with experiments, see for instance: ZIMMERMANN199199 ; PhysRevB.54.700 ; Dressel2013-mq ; Michon2023-lq ; Boyack2023-yu .
On the other hand, regarding the imaginary part $\text{Im}[\sigma]$, the functional dependence on the frequency is simpler as it is defined by the function $\mathcal{G}(a,\omega)$ in 38b. The low-frequency behavior is given by $\text{Im}[\sigma]\sim 1/\omega$, while there is a minimum above which the behavior follows a linear trend, $\text{Im}[\sigma]\sim\omega$. This holds even in the scenario without rotation. Nonetheless, as seen in Figure 3(b), notable distinctions emerge: in the scenario devoid of rotation $a=0$, the minimum is located at a significantly higher frequency compared to cases with nontrivial rotation. Additionally, the slope of the dominant linear function at high frequencies is smaller, leading to a slower-growing behavior. To end this section, let us remark that our analytical findings (solid lines) show strong correspondence with the numeric (dashed lines) as occurred in the previous scenario. Nevertheless, regarding $\text{Re}[\sigma]$, the correspondence is sufficiently good only when the rotation parameter is small, for example within the interval $40\leq a\leq 1/$ 4. For higher values of the rotation parameter, the analytical curves of $\text{Re}[\sigma]$ shift to the right compared to the numerical ones, see Figure 4.
Figure 4: Illustration of the shift in $\text{Re}[\sigma]$ for different values of the rotation parameter. The analytical curves are shifted to the right (solid) with respect to the numerical ones (dashed) as the rotation parameter adopts higher values. For this plot, we choose $T\approx 0.99994\,T_{c}$ for which $\expectationvalue{\mathcal{O}}\approx 7.780$, with parameters $\ell=1$ and $r_{h}=1$.

### v2

@@ 4 The AC conductivity
The AC conductivity is a fundamental quantity used to characterize a superconductor, as it captures the response of the system to external electric perturbations. In our holographic setup, this can be analyzed by introducing a gauge field perturbation within the bulk dynamics and solving the resulting linearized Maxwell equations. Specifically, we consider a perturbation in the $x$-component of the gauge field, $\delta A_{x}$, with a harmonic time dependence: $\delta A_{x}=A_{x}(u)e^{-i\omega t}$. Thus, after substituting this ansatz into (7), we obtain
@@
To solve this equation, we impose ingoing wave boundary conditions at the horizon
@@
while the behavior at the asymptotic boundary is given by
@@
where, from the standard holographic dictionary, we interpret $A_{x}^{(0)}$ as the source of a conjugated current $J_{x}$, while $A_{x}^{(z+1)}$ is related to its expectation value, $\expectationvalue{J_{x}}\sim A_{x}^{(z+1)}$. Applying the Ohm’s law, we thus obtain an expression for the AC conductivity
@@
In the remainder of this section, we focus in analyzing the frequency and angular momentum dependence of the conductivity. For this purpose, we conduct an analytical study of the Maxwell perturbation (16) considering the near-boundary limit of the background fields. We carry out this analysis by approximating $f\approx 1$ and $\psi\approx\expectationvalue{\mathcal{O}}u^{\Delta}/\sqrt{2}$, yielding
@@
Analytical solutions to this equation are only obtainable for specific values of the parameters $\Delta$ and $z$. In what follows, we consider the case $\Delta=1$, which corresponds to a scalar mass squared of $m^{2}=-(z+2)/\ell^{2}$. This value is consistent with the BF bound of Lifshitz spacetimes, ensuring the stability of the scalar sector. We present two specific examples where equation (20) admits analytic solutions. These results are further supported by numerical computations, in which we solve the full equation (16) subject to the boundary conditions (17) and (18).
4.1 Isotropic case: $\Delta=1$ and $z=1$
In this case, the equation (20) admits two independent solutions, which can be expressed in terms of Bessel functions of the first, $J_{1}$, and second kind, $Y_{1}$. Namely,
@@ 4.1 Isotropic case: $\Delta=1$ and $z=1$
Before proceeding, it is convenient to impose the following relation between the integration constants
$\dfrac{c_{0}}{c_{1}}=\dfrac{i}{\pi}\left[\pi H(\omega-\omega_{g})+(1-2\gamma_{E})i-\dfrac{180}{h^{2}(\omega)}+\frac{i}{2}\,\ln\left(\dfrac{h^{2}(\omega)}{4}\right)-\pi\right],$ (23)
@@
in which $H(\omega-\omega_{g})$ denotes the Heaviside step function, and $\gamma_{E}$ is the Euler-Mascheroni constant. The asymptotic analysis of the solution (21), combined with the relation (19), yields the following expression for the conductivity
$\sigma(\omega)=-\frac{{h^{2}(\omega)}}{\omega}\left(\pi H(\omega-\omega_{g})+\dfrac{i}{2}\ln\left(\dfrac{h^{2}(\omega)}{4}\right)-\,\dfrac{180\,i}{h^{2}(\omega)}\right).$ (24)
@@
A closer inspection of this relation uncovers the structure of real and imaginary parts of the conductivity, leading to
[…]
@@
Here, $\omega_{g}=\expectationvalue{\mathcal{O}}/\ell^{2}$ denotes a gap frequency determined by the function $h(\omega)$ below which the real part of the conductivity vanishes—except at $\omega=0$, where a delta function appears.
@@
In Figure 2, we show the behavior of the real and imaginary parts of the conductivity defined in the previous relations (25)-(26). The observed features exhibits recurring characteristics of those reported in previous holographic superconductor models. Regarding the real part of the conductivity, the presence of a delta function at $\omega=0$, indicates the emergence of a DC current and signals the onset of the superconducting state. Moreover, the existence of a gap frequency $\omega_{g}$, proportional to the order parameter $\expectationvalue{\mathcal{O}}$, is also a common property in holographic superconductors. However, it has been argued that holographic superconductors are not truly gapped; instead, the apparent gap arises from decoupling the matter sector from gravity, a consequence of employing the probe limit approximation. For an in-depth discussion, we refer the reader to Hartnoll:2009sz ; Horowitz:2009ij ; Horowitz:2010gk . For frequencies above the gap frequency $\omega>\omega_{g}$, the real part of the conductivity follows a monotonic increase approximately given by $\text{Re}[\sigma]\sim-h^{2}(\omega)/\omega$, following a linear tendency $\text{Re}[\sigma]\sim\omega$ at sufficiently higher frequencies.
@@
Turning our attention to the imaginary component (26), we observe a minimum occurring very close to $\omega_{g}$. Below this minimum, the imaginary part exhibits a sharp increase characterized by $\text{Im}[\sigma]\sim 1/\omega$. Conversely, at sufficiently higher frequencies $\omega\gg\omega_{g}$, the behavior is given by $\text{Im}[\sigma]\sim\omega\ln\left(\ell^{2}\omega/2r_{h}\right)$. These features are illustrated in Figure 2, where we highlight the excellent agreement between our analytical predictions (solid lines) and numerical results (dashed lines).
@@
Lastly, it is important to note that we have not addressed the rotation of the background in the previous scenario. The reason for this is that neither of the expressions, (25) or (26), possess information on the rotation parameter $a$. This observation results evident from the effective equation (20), where the absence of anisotropy ($z=1$) leads to the cancellation of terms associated with rotation, rendering no residual information on this parameter. Consequently, we infer that anisotropy ($z\neq 1$) is an essential ingredient to capture the impact of rotation on the AC conductivity, at least within the scope of our holographic model.
@@
Figure 2: Real (left) and imaginary (right) parts of the conductivity as functions of the frequency. The solid curves represent the analytical results, while the dashed ones correspond to numerical computations. These plots were generated at $T\approx 0.2\,T_{c}$, with parameters $\Delta=1$, $z=1$, $\ell=1$ and $r_{h}=1$.
4.2 Anisotropic case: $\Delta=1$ and $z=2$
@@
In this scenario, the solution can be written in terms of the confluent hypergeometric functions of the first kind, $M$, and second kind, $U$,
@@
As in the preceding case, we can obtain an analytic expression for the conductivity by employing the relation (19). To accomplish this, we find it convenient to introduce the following identification
@@
$\mathcal{F}(a)$ $=\frac{711}{400}+\quantity(\dfrac{a}{\ell})\exp[-\quantity(\dfrac{a}{\ell})^{2}]\Biggl{(}\frac{3}{32}-\frac{464}{249}\quantity(\dfrac{a}{\ell})+\frac{829}{100}\quantity(\dfrac{a}{\ell})^{2}-\frac{22721}{1000}\quantity(\dfrac{a}{\ell})^{3}$
$\quad+\frac{6497}{250}\quantity(\dfrac{a}{\ell})^{4}\Biggr{)},$ (30a)
$\mathcal{G}(a,\omega)$ $=3\ell\Biggl{[}6.61\times 10^{4}\left(\dfrac{a}{\ell}\right)^{4}+1.17\times 10^{4}\left(\dfrac{a}{\ell}\right)^{2}-763\left(\dfrac{a}{\ell}\right)+\dfrac{3}{2}+\dfrac{6.61\times 10^{6}}{\ell\omega}\left(\dfrac{\expectationvalue{\mathcal{O}}}{\ell}\right)^{2}$
$\quad+3.33\times 10^{6}\left(\dfrac{a}{\ell}\right)^{2}\ell\omega\Biggr{]}.$ (30b)
Altogether, incorporating the definitions above, we arrive at the following analytic expression for the conductivity
@@
[…]
@@
When the rotation is switched on ($a\neq 0$), the asymptotic behavior of $\text{Re}[\sigma]$ undergoes a notable transformation. In this scenario, the Dirac delta function at $\omega=0$ remains, but it is now accompanied by the appearance of a narrow gap at low frequencies. This discernible gap can be characterized by noting that the values assumed by $\text{Re}[\sigma]$ within the interval $0<\omega\leq\omega_{g}$ are significantly smaller than those beyond this range, $\omega>\omega_{g}$. As a result, within the region $0<\omega\leq\omega_{g}$, we can approximate $\text{Re}[\sigma]$ effectively as zero. Under this assumption, we can estimate the width of the gap as a function of the rotation parameter, leading to
$\omega_{g}/T_{c}\approx 8\quantity(\dfrac{a}{\ell})^{2}+10.5,$ (32)
which fits very well for small values of the rotation, particularly within $0\leq a\leq 2/5$. Moreover, this equation indicates that the gap persists even in the absence of rotation, as expected from our earlier discussion. We note that the increase of the frequency gap with the rotation parameter indicates that, once the condensate is formed, the superconducting state becomes more robust against excitations, as a larger amount of energy is required to break the Cooper pairs. In this sense, the system departs from the fundamental BCS prediction ($\omega_{g}/T_{c}\approx 3.5$) and moves into a strong–coupling regime reminiscent of high-$T_{c}$ superconductivity.
@@
Figure 3: Portraits of the real (left) and imaginary (right) components of the conductivity as a function of the frequency. Note that the analytic curves (solid lines) show an excellent agreement with numerical ones (dashed lines). The plots were generated at $T\approx 0.061\,T_{c}$, considering $\Delta=2$, $z=2$, $\ell=1$ and $r_{h}=1$.
It is interesting to notice that, when $\omega>\omega_{g}$, with the rotation switched on ($a\neq 0$), the curve corresponding to $\text{Re}[\sigma]$ exhibits a peak, beyond which it undergoes an exponential fall. In this high frequency regime, the term $a^{2}\omega/4\ell\Xi$ dominates over $\expectationvalue{\mathcal{O}}/\ell\omega$ in the function $g(\omega,a)$, leading to an asymptotic behavior of the form
$\text{Re}[\sigma]\sim\omega^{2}\exp\left(-2\pi a^{2}\omega/4\ell\Xi\right),$ (33)
which explicitly reflects the dependence on the rotation parameter. Therefore, the combined effect of rotation and anisotropy in the background is accountable for inducing the peak and vanishing exponential behavior. This is a noteworthy characteristic in our model, as it closely resembles the behavior reported in both theoretical and experimental studies of high-$T_{c}$ superconductivity (see, for instance: ZIMMERMANN199199 ; PhysRevB.54.700 ; Dressel2013-mq ; Michon2023-lq ; Boyack2023-yu ).
On the other hand, regarding the imaginary part $\text{Im}[\sigma]$, the functional dependence on the frequency is simpler as it is defined by the function $\mathcal{G}(a,\omega)$ in 30b. At low frequencies, it follows the inverse behavior $\text{Im}[\sigma]\sim 1/\omega$, while beyond a certain minimum —occurring at
$\omega_{\text{min}}\approx\dfrac{2\expectationvalue{\mathcal{O}}}{\ell a},$ (34)
for $a\neq 0$, it transits into a linear regime characterized by $\text{Im}[\sigma]\sim\omega$. Nonetheless, as expected from equation (30b), such a minimum does not appear in the non-rotating case ($a=0$). In this case, the function $\text{Im}[\sigma]$ approaches to a constant in the high frequency limit, as shown by the black curve in the right panel of Figure 3.
Regarding the present analysis, note that our analytical results (solid lines) show excellent agreement with the numerical computations (dashed lines). However, this correspondence remains accurate primarily for small values of the rotation parameter, typically within the interval $0\leq a\leq 2/5$. This is because the functions $\mathcal{F}(a)$ and $\mathcal{G}(a,\omega)$ were specifically chosen to match the analytic and numerical results within this range of values for $a$.
To conclude this section, let us examine the temperature dependence of the conductivity. From Figure 4, we observe that as the temperature increases, the real part of the conductivity shifts upward, while the imaginary part shifts downward. Moreover, the apparent gap in $\text{Re}[\sigma]$ observed at low temperatures gradually closes and eventually vanishes at high temperatures. A plausible suggestion is that the estimated gap (32) is, in general, a function of both temperature and rotation, since it is determined by the condensate $\expectationvalue{\mathcal{O}}$, which itself varies with these parameters.
Figure 4: Real and imaginary parts of the conductivity as a function of the frequency for different temperatures. For this plot, we choose $a=1/20$, $\ell=1$ and $r_{h}=1$.